# Revealing the building blocks of tree balance: fundamental units of the Sackin and Colless Indices

### v1

@@ 3.1.3 Other properties of $\Delta_{CS}$, $\Delta_{SC}$ and $N_{b}$
  • $\lambda=(0,1)$ as base case for trees with $n=1$ leaf,
  • for trees $T=(T_{a},T_{b})$ we have recursions $r_{1}(T_{a},T_{b})=\Delta_{CS}(T_{a})+\Delta_{CS}(T_{b})+|n_{a}-n_{b}|-(n_{a}+n_{b})$ ($\Delta_{CS}$ index) and $r_{2}(T_{a},T_{b})=n_{a}+n_{b}$ (leaf numbers).
@@ 3.2 $N_{a}$: another new imbalance index on $\mathcal{BT}^{*}_{n}$
Seeking a contradiction, suppose there is a tree $T=(T_{a},T_{b})$ minimizing $N_{a}$ but with $T_{a}$ not minimizing $N_{a}$ (the proof in case $T_{b}$ is not minimal follows analogously). This implies that there is a tree $T_{a}^{\prime}\in\mathcal{BT}^{*}_{n_{a}}$ such that $N_{a}(T_{a}^{\prime})<N_{a}(T_{a})$. We now consider $T^{\prime}=(T_{a}^{\prime},T_{b})$, i.e., $T^{\prime}$ is like $T$ except that $T_{a}$ has been replaced by $T_{a}^{\prime}$. Using Lemma 7, this leads to:
$N_{a}(T)$ $=N_{a}(T_{a})+N_{a}(T_{b})+n_{a}>N_{a}(T_{a}^{\prime})+N_{a}(T_{b})+n_{a}=N_{a}(T^{\prime}),$
contradicting the $N_{a}$ minimality of $T$. This completes the proof. ∎
@@
We first consider the Sackin index and recall that $S=N_{a}+N_{b}$, and thus $N_{a}=S-N_{b}=S+(-N_{b})$. Furthermore, recall that if $T$ is a Colless minimum, it is also a Sackin minimum (cf. Section 2.2.4) as well as an $N_{b}$ maximum by Theorem 3 and thus a $N_{b}$ minimum. Therefore, if $T$ is a Colless minimum, both summands in $N_{a}=S+(-N_{b})$ are minimal, which shows that indeed every Colless minimum also minimizes $N_{a}$.
For the other direction, seeking a contradiction, we assume $T=(T_{a},T_{b})$ minimizes $N_{a}$ but is not a Colless minimum. Without loss of generality we may assume $T$ is minimal with this property; i.e., all smaller $N_{a}$ minimal trees are also Colless minimal. Then we distinguish two cases based on $n_{a}$ and $n_{b}$ – the numbers of leaves of $T_{a}$ and $T_{b}$ with $n_{a}\geq n_{b}$:
@@
    If $n_{a}-n_{b}\leq n_{a}^{gfb}-n_{b}^{gfb}$, we know that $T$ is Sackin minimal by Section 2.2.3. By Corollary 3 we know that $T_{a}$ and $T_{b}$ also minimize $N_{a}$ (as $T$ does so), but since they have fewer leaves than $T$, by the minimal choice of $T$ we know that $T_{a}$ and $T_{b}$ are also Colless minimal. Thus, the only way we can have $n_{a}-n_{b}\leq n_{a}^{gfb}-n_{b}^{gfb}$ without $T=(T_{a},T_{b})$ being a Colless minimum is – as both maximal pending subtrees $T_{a}$ and $T_{b}$ are Colless minima – that $(n_{a},n_{b})$ is not a Colless partition. In particular, it cannot be the gfb-partition $(n_{a}^{gfb},n_{b}^{gfb})$.
    We now argue that it cannot be the echelon partition either. Let $k_{n}=\lceil\log_{2}n\rceil$. If $n\geq 3\cdot 2^{k_{n}-2}$, we have $n_{a}^{gfb}=2^{k_{n}-1}=n_{a}^{be}$ by Section 2.2.1, implying that in this case, the echelon partition coincides with the gfb-partition, which we have already excluded.
    If, however, $n<3\cdot 2^{k_{n}-2}$, we have $n_{a}^{gfb}-n_{b}^{gfb}=(n-2^{k_{n}-2})-2^{k_{n}-2}=n-2^{k_{n}-1}<3\cdot 2^{k_{n}-2}-2^{k_{n}-1}=2^{k_{n}-2}$ (cf. Section 2.2.1). Moreover, we have: $n_{a}^{be}-n_{b}^{be}=2^{k_{n}-1}-(n-2^{k_{n}-1})=2^{k_{n}}-n>2^{k_{n}}-3\cdot 2^{k_{n}-2}=2^{k_{n}-2}$. This shows that $n_{a}-n_{b}\leq n_{a}^{gfb}-n_{b}^{gfb}=2^{k_{n}-2}<n_{a}^{be}+n_{b}^{be}$, implying that $(n_{a},n_{b})$ cannot be the echelon partition in this case, either.
    As $(n_{a},n_{b})$ is neither a Colless partition nor the echelon partition, by Theorem 5, $T$ cannot maximize $N_{b}$ and thus, in turn, not minimize $N_{b}$. Now, let $T^{\prime}$ be a Colless minimum (which also implies Sackin minimality by Section 2.2.4). Then we know by Section 2.2.4 that $T^{\prime}$ is also a Sackin minimum, and we furthermore know, by Theorem 3, that $T^{\prime}$ maximizes $N_{b}$ and thus minimizes $N_{b}$. Thus, we get:
@@
    Let $T^{\prime}$ be a Colless minimum. We then know, by Section 2.2.4, that $T^{\prime}$ is also a Sackin minimum, and we also know, by Theorem 3, that $T^{\prime}$ maximizes $N_{b}$ and thus minimizes $N_{b}$. Thus, we get:
@@
Consider $T$ to be an ordered tree such that for each inner vertex $v$ with children $v_{a}$ and $v_{b}$, whenever $n_{v_{a}}\geq n_{v_{b}}$, we have that $v_{a}$ is the “left” child of $v$ and $v_{b}$ is the “right” child of $v$. This way, each edge in $T$ can be labeled with $L$ or $R$, depending on whether it is leading to the left or the right child of its tail vertex.
Now, if you consider a leaf $x\in V^{1}$, this implies that the unique path $P_{x}$ from the root $\rho$ of $T$ to $x$, which defines the depth $\delta_{x}$, can be subdivided into $L$-edges and $R$-edges. We denote by $\delta_{x}^{L}$ and $\delta_{x}^{R}$ the number of $L$- and $R$-edges on this path, respectively. Then, we get:
@@
$N_{b}=\sum\limits_{v\in\mathring{V}(T)}n_{v_{a}}=\sum\limits_{x\in V^{1}(T)}\delta^{R}_{x}.$
@@
Next, we show that $N_{a}$ is local.
@@
Proof.
Let $T$, $T^{\prime}\in\mathcal{BT}^{*}_{n}$ with $T^{\prime}$ obtained from $T$ by replacing the pending subtree $T_{v}$ by a subtree $T_{v}^{\prime}$ of the same size. Using the identity $N_{a}(T)=S(T)-N_{b}(T)$ as well as the locality of Sackin and $N_{b}$ (cf. Proposition 7), we obtain
$N_{a}(T)-N_{a}(T^{\prime})$ $=(S(T)-N_{b}(T))-(S(T^{\prime})-N_{b}(T^{\prime}))$
$=(S(T)-S(T^{\prime}))-(N_{b}(T)-N_{b}(T^{\prime}))$
$=(S(T_{v})-S(T_{v}^{\prime}))-(N_{b}(T_{v})-N_{b}(T_{v}^{\prime}))$
$=(S(T_{v})-N_{b}(T_{v}))-(S(T_{v}^{\prime})-N_{b}(T_{v}^{\prime}))$
$=N_{a}(T_{v})-N_{a}(T_{v}^{\prime}).$
@@
  • for trees $T=(T_{a},T_{b})$ we have recursions $r_{1}(T_{a},T_{b})=N_{a}(T_{a})+N_{a}(T_{b})+\frac{1}{2}(|n_{a}-n_{b}|+(n_{a}+n_{b}))$ ($N_{a}$ index) and $r_{2}(T_{a},T_{b})=n_{a}+n_{b}$ (leaf numbers).
@@
We are now in a position to present several new results on the echelon tree.
@@ 3.3 Additional results on echelon trees $T^{be}_{n}$
Now, let again $k_{n}=\lceil\log_{2}(n)\rceil$ and consider the binary expansion of $n$, i.e., $n=\sum\limits_{i=0}^{k_{n}}\alpha_{i}\cdot 2^{i}$, where $\alpha_{i}$ equals 1 precisely if $2^{i}$ is contained in the binary expansion of $n$ and 0 else. Note that $w(n)=\sum\limits_{i=0}^{k_{n}}\alpha_{i}$ is the so-called binary weight of $n$. Let $f_{n}(i)$ for $i=1,\ldots,w(n)$ denote the $i^{th}$ power of two in the binary expansion of $n$ when they are sorted in ascending order. For instance, if $n=11=2^{0}+2^{1}+2^{3}$, we have $f_{11}(1)=2^{0}$, $f_{11}(2)=2^{1}$ and $f_{11}(3)=2^{3}$. Note that this implies $n=\sum\limits_{i=1}^{w(n)}f_{n}(i)$.
@@
Next, we consider the case in which $n$ is not a power of two and prove the statement by induction on $n$, starting with $n=3$. In the following, let $k_{n}=\lceil\log_{2}(n)\rceil$. For the base case $n=3=2^{1}+2^{0}$ we have $w(3)=2$ with $f_{3}(1)=2^{0}$ and $f_{3}(2)=2^{1}$. Hence, the top caterpillar of $T^{*}_{3}$ is $T^{cat}_{2}$ and the trees that replace its two leaves are $T^{fb}_{\log_{2}(f_{3}(1))}=T^{fb}_{0}$ and $T^{fb}_{\log_{2}(f_{3}(2))}=T^{fb}_{1}$, resulting in $T^{*}_{3}=T^{cat}_{3}$. On the other hand, $k_{3}=2$, and by the recursive definition of the echelon tree, $T_{3}^{be}$ consists of the maximal pending subtrees $T^{fb}_{k_{n}-1}=T^{fb}_{2-1}=T_{1}^{fb}$ and $T^{be}_{n-2^{k_{n}-1}}=T^{be}_{3-2^{2-1}}=T^{be}_{1}$. Since, by definition, $T^{be}_{1}$ is a single vertex, we have $T^{be}_{n}=T^{cat}_{3}=T^{*}_{3}$, completing the base case.
@@
We conclude this section by providing explicit formulas for the various different indices discussed in our manuscript for the echelon tree, which are new to the literature.
@@
    $\sum\limits_{v\in\mathring{V}(T^{cat}_{w(n)})}n_{v}$ $=\sum\limits_{i=1}^{w(n)-1}n_{v_{i}}$
    $=\sum\limits_{i=1}^{w(n)-1}\left(f_{n}(i+1)+\sum\limits_{j=1}^{i}f_{n}(j)\right)$
    $=\left(\sum\limits_{i=1}^{w(n)-1}f_{n}(i+1)\right)+f_{n}(1)+(f_{n}(1)+f_{n}(2))+\ldots+(f_{n}(1)+\ldots+f_{n}(w(n)-1))$
    $=\left(\sum\limits_{i=1}^{w(n)-1}f_{n}(i+1)\right)+\underbrace{f_{n}(1)+\ldots+f_{n}(1)}_{w(n)-1\ \text{times}}+\underbrace{f_{n}(2)+\ldots+f_{n}(2)}_{w(n)-2\ \text{times}}+\ldots+\underbrace{f_{n}(w(n)-1)}_{1\ \text{time}}$
@@
    This explains the last summand of the term and thus completes the proof for the Sackin index.
  • For the Colless index, we proceed as for the Sackin index. We note, however, that each fully balanced subtree contributes a Colless index of 0. Hence, the respective term that, for the Sackin index led to the second sum, is 0 in the Colless case and thus does not need to be considered.
@@
  • Finally, as we know that the Sackin index equals $S=N_{a}+N_{b}$, we have $N_{a}=S-N_{b}$. The formula for $N_{a}$ is now a direct conclusion of the already derived formulas for $S$ and $N_{b}$. This completes the proof.
@@
Note that the integer sequences 0,2,5,8,13,16,20,24,33,36,40 and 0,0,1,0,3,2,2,0,7,6,6,4,6,4,3,0 and 0,-2,-4,-8,-10,-14,-18,-24,-26,-30 and 0,2,4,8,10,14,18,24,26,30 of the values of $S(T_{n}^{be})$, $C(T_{n}^{be})$, $\Delta_{CS}$ and $\Delta_{SC}$, respectively, for $n\geq 1$ had not been contained in the Online Encyclopedia of Integer sequences [24]. This indicates that they might not have occurred in other contexts yet. They have been added to the OEIS in the course of this manuscript. Interestingly, the sequences 0,1,3,4,8,9,11,12,20,21,23,24 of $N_{a}(T_{n}^{be})$ had, at least without the first term of 0 for the trivial tree consisting of only one vertex, already be contained in the OEIS [24, Sequence A006520]. Also, the sequence 1,2,4,5,7,9,12,13,15,17,20,22,25,28,32,33,35, 37,40,42,45 of $N_{b}(T_{n}^{be})$ had already been contained in the OEIS [24, Sequence A000788]. The fact that these sequences have already occurred in other contexts nicely links our findings on the echelon tree and the new indices $N_{a}$ and $N_{b}$ to other combinatorial research areas.

### v2

@@ 3.1.3 Further properties of $\Delta_{CS}$, $\Delta_{SC}$ and $N_{b}$
  • $\lambda=(0,1)$ as base case for the tree with $n=1$ leaf,
  • for trees $T=(T_{a},T_{b})$ we have recursions $r_{1}(T_{a},T_{b})=\Delta_{CS}(T_{a})+\Delta_{CS}(T_{b})+|n_{a}-n_{b}|-(n_{a}+n_{b})$ ($\Delta_{CS}$ value) and $r_{2}(T_{a},T_{b})=n_{a}+n_{b}$ (leaf numbers).
@@ 3.2 $N_{a}$: another new imbalance index on $\mathcal{BT}^{*}_{n}$
The proof of Corollary 3 is analogous to the proof of Corollary 1, except that Lemma 8 is used instead of Proposition 1. ∎
@@
First, recall that if $T$ is a Colless minimum, it is also a Sackin minimum (cf. Section 2.2.4) as well as an $N_{b}$ minimum by Theorem 3. Therefore, if $T$ is a Colless minimum, both summands in $N_{a}=S+(-N_{b})$ are minimal, which shows that indeed every Colless minimum also minimizes $N_{a}$.
For the other direction, seeking a contradiction, we assume $T=(T_{a},T_{b})$ minimizes $N_{a}$ but is not a Colless minimum. Without loss of generality we may assume $T$ is minimal with this property. Then we distinguish two cases based on $n_{a}$ and $n_{b}$ – the numbers of leaves of $T_{a}$ and $T_{b}$ with $n_{a}\geq n_{b}$:
@@
    First, consider the case $n_{a}-n_{b}\leq n_{a}^{gfb}-n_{b}^{gfb}$. As $T$ minimizes $N_{a}$, so do $T_{a}$ and $T_{b}$, and by the minimality of $T$, $T_{a}$ and $T_{b}$ are Colless minimal. This implies that $(n_{a},n_{b})$ cannot be a Colless partition (else $T$ would be a Colless minimum), which in particular shows $(n_{a},n_{b})\neq(n_{a}^{gfb},n_{b}^{gfb})$. Thus, it cannot be the gfb partition.
    We now argue that it cannot be the echelon partition either. Let $k_{n}=\lceil\log_{2}n\rceil$. If $n\geq 3\cdot 2^{k_{n}-2}$, we have already seen that in this subcase the echelon partition coincides with the gfb-partition, which we have already excluded.
    If, however, $n<3\cdot 2^{k_{n}-2}$, just as in the second part of the proof of Proposition 4 we have $n_{a}-n_{b}<n_{a}^{gfb}-n_{b}^{gfb}<n_{a}^{be}-n_{b}^{be}$, implying that $(n_{a},n_{b})$ cannot be the echelon partition in this case, either.
    As $(n_{a},n_{b})$ is neither a Colless partition nor the echelon partition, by Theorem 6, $T$ cannot minimize $N_{b}$. Now, let $T^{\prime}$ be a Colless minimum (which also implies Sackin minimality). Then we know by Theorem 3, that $T^{\prime}$ minimizes $N_{b}$. Thus, we get:
@@
    Let $T^{\prime}$ be a Colless minimum. As above, $T^{\prime}$ then minimizes $S$ and $N_{b}$, too. Thus, we get:
@@
Consider $T$ to be an ordered tree such that for each inner vertex $v$ with children $v_{a}$ and $v_{b}$, whenever $n_{v_{a}}\geq n_{v_{b}}$, we have that $v_{a}$ is the “left” child of $v$ and $v_{b}$ is the “right” child of $v$. In this way, each edge in $T$ can be labeled with $L$ or $R$, depending on whether it is leading to the left or the right child of its tail vertex.
Now, if we consider a leaf $x\in V^{1}$, this implies that the unique path $P_{x}$ from the root $\rho$ of $T$ to $x$, which defines the depth $\delta_{x}$, can be subdivided into $L$-edges and $R$-edges. We denote by $\delta_{x}^{L}$ and $\delta_{x}^{R}$ the number of $L$- and $R$-edges on this path, respectively. Then, we get:
@@
$N_{b}=\sum\limits_{v\in\mathring{V}(T)}n_{v_{b}}=\sum\limits_{x\in V^{1}(T)}\delta^{R}_{x}.$
@@
Next, we state that $N_{a}$ is local.
@@
We omit the proof as it is completely analogous to that of Proposition 7, except that $N_{a}$ now plays the role of $\Delta_{CS}$, $S$ plays the role of $C$ and $N_{b}$ plays the role of $S$.
@@
  • for trees $T=(T_{a},T_{b})$ we have recursions $r_{1}(T_{a},T_{b})=N_{a}(T_{a})+N_{a}(T_{b})+\frac{1}{2}(|n_{a}-n_{b}|+(n_{a}+n_{b}))$ ($N_{a}$ value) and $r_{2}(T_{a},T_{b})=n_{a}+n_{b}$ (leaf numbers).
@@
Since the echelon tree plays an important role in our manuscript and as the existing literature on it is very sparse, we will now present several new results on the echelon tree.
@@ 3.3 Additional results on echelon trees $T^{be}_{n}$
Now, let again $k_{n}=\lceil\log_{2}(n)\rceil$ and consider the binary expansion of $n$, i.e., $n=\sum\limits_{i=0}^{k_{n}}\alpha_{i}\cdot 2^{i}$, where $\alpha_{i}$ is the $i+1$-th digit of the binary representation of $n$. Note that $w(n)=\sum\limits_{i=0}^{k_{n}}\alpha_{i}$ is the so-called binary weight of $n$. Let $f_{n}(i)$ for $i=1,\ldots,w(n)$ denote the $i^{th}$ power of two in the binary expansion of $n$ when they are sorted in ascending order. For instance, if $n=11=2^{0}+2^{1}+2^{3}$, we have $f_{11}(1)=2^{0}$, $f_{11}(2)=2^{1}$ and $f_{11}(3)=2^{3}$. Note that this implies $n=\sum\limits_{i=1}^{w(n)}f_{n}(i)$.
@@
Next, we consider the case in which $n$ is not a power of two and prove the statement by induction on $n$, starting with $n=3$. In the following, let $k_{n}=\lceil\log_{2}(n)\rceil$. For the base case $n=3=2^{1}+2^{0}$ we have $w(3)=2$ with $f_{3}(1)=2^{0}$ and $f_{3}(2)=2^{1}$. Hence, the top caterpillar of $T^{*}_{3}$ is $T^{cat}_{2}$ and the trees that replace its two leaves are $T^{fb}_{\log_{2}(f_{3}(1))}=T^{fb}_{0}$ and $T^{fb}_{\log_{2}(f_{3}(2))}=T^{fb}_{1}$, resulting in $T^{*}_{3}=T^{cat}_{3}$. On the other hand, $k_{3}=2$, and by the recursive definition of the echelon tree, $T_{3}^{be}$ consists of the maximal pending subtrees $T^{fb}_{k_{n}-1}=T^{fb}_{2-1}=T_{1}^{fb}$ and $T^{be}_{n-2^{k_{n}-1}}=T^{be}_{3-2^{2-1}}=T^{be}_{1}$. Since, by definition, $T^{be}_{1}$ is a single vertex, we have $T^{be}_{3}=T^{cat}_{3}=T^{*}_{3}$, completing the base case.
@@
We conclude this section by providing explicit formulas for the various indices discussed in our manuscript for the echelon tree, which are new to the literature.
@@
    $\sum\limits_{v\in\mathring{V}(T^{cat}_{w(n)})}n_{v}$ $=\sum\limits_{i=1}^{w(n)-1}\underbrace{\left(f_{n}(i+1)+\sum\limits_{j=1}^{i}f_{n}(j)\right)}_{n_{v_{i}}}$
    $=\left(\sum\limits_{i=1}^{w(n)-1}f_{n}(i+1)\right)+\left(\sum\limits_{i=1}^{w(n)-1}\sum\limits_{j=1}^{i}f_{n}(j)\right)$
@@
    This explains the first summand of the term and thus completes the proof for the Sackin index.
  • For the Colless index, we proceed as for the Sackin index. We note, however, that each fully balanced subtree contributes a Colless index of 0. Hence, the respective term that for the Sackin index led to the second sum, is 0 in the Colless case and thus does not need to be considered.
@@
  • Finally, as we know that $N_{a}=S-N_{b}$, the formula for $N_{a}$ is a direct conclusion of the already derived formulas for $S$ and $N_{b}$. This completes the proof.
@@
Note that the integer sequences 0,2,5,8,13,16,20,24,33,36,40 and 0,0,1,0,3,2,2,0,7,6,6,4,6,4,3,0 and 0,-2,-4,-8,-10,-14,-18,-24,-26,-30 and 0,2,4,8,10,14,18,24,26,30 of the values of $S(T_{n}^{be})$, $C(T_{n}^{be})$, $\Delta_{CS}(T_{n}^{be})$ and $\Delta_{SC}(T_{n}^{be})$, respectively, for $n\geq 1$ had not been contained in the Online Encyclopedia of Integer sequences [27]. This indicates that they might not have occurred in other contexts yet. They have been added to the OEIS in the course of this manuscript. Interestingly, the sequences 0,1,3,4,8,9,11,12,20,21,23,24 of $N_{a}(T_{n}^{be})$ had, at least without the first term of 0 for the trivial tree consisting of only one vertex, already been contained in the OEIS [27, Sequence A006520]. Also, the sequence 1,2,4,5,7,9,12,13,15,17,20,22,25,28,32,33,35, 37,40,42,45 of $N_{b}(T_{n}^{be})$ had already been contained in the OEIS [27, Sequence A000788]. The fact that these sequences have already occurred in other contexts nicely links our findings on the echelon tree and the new indices $N_{a}$ and $N_{b}$ to other combinatorial research areas.